cash now rde tt utre duia hom the folo ng finanaal sta monts in tho popup window Tho company pad trott exporoo o $17,900 or 201アand had an overall tax rate of 40% to 2017. erty the cat fo do th The cash fow from assets is$(Round to the nearest dolar) Data Table $348 Cost of $141 542.000 $27,900 $45 Shoot 12/31/2016 TS LIABILITIES 15,9D0 Ncles payable 27,800 Accounts payable $47,B00|Lang-larm dett 189 S368,000OWNERS EQUITY 142,500 Retained 8amings 82,200 Common slock $131 Sheet 12/31/2017 LIABILITIES 26,000Nctos payablo 18,800 Accounts payable $52,800 Long torm dobt rocovable 448,000OWNERS EQUITY d eamings nd assets 82,200 Cemman slack $181 Enter your anewer in the arewer box and then cick Check Anewer. Frint Done Check Answer
Answer #1

Cash Flow to Creditors = Interest - New Debt = 17900 - ( 161900 - 189800) = 45800
Cash flow to owners = -Increase in Retained Earnings and Common Equity =-(181,900-131,800) =-51,100
Cash Flow from Assets = 45800 -51,100 = -4300

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well